OpenStudy (anonymous):

Complex number √-6(2+√-8)

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{-6}(2+\sqrt{-8})\] \[\Large \sqrt{-6}*2+ \sqrt{-6}*\sqrt{-8}\] \[\Large 2\sqrt{-6}+ \sqrt{-6}*\sqrt{-8}\] \[\Large 2\sqrt{-1}*\sqrt{6}+ \sqrt{-1}*\sqrt{6}*\sqrt{-1}*\sqrt{8}\] Keep going to simplify

OpenStudy (anonymous):

There's the thing, if I could simplify I would... I got up to there and then stopped...

jimthompson5910 (jim_thompson5910):

keep in mind that \[\Large i = \sqrt{-1}\]

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{-6}(2+\sqrt{-8})\] \[\Large \sqrt{-6}*2+ \sqrt{-6}*\sqrt{-8}\] \[\Large 2\sqrt{-6}+ \sqrt{-6}*\sqrt{-8}\] \[\Large 2\sqrt{-1}*\sqrt{6}+ \sqrt{-1}*\sqrt{6}*\sqrt{-1}*\sqrt{8}\] \[\Large 2i*\sqrt{6}+ i*\sqrt{6}*i*\sqrt{8}\] \[\Large 2i*\sqrt{6}+ i*i*\sqrt{6}*\sqrt{8}\] \[\Large 2i*\sqrt{6}+ i^2*\sqrt{6*8}\] \[\Large 2i*\sqrt{6}+ (-1)*\sqrt{48}\] \[\Large 2i*\sqrt{6}+ (-1)*\sqrt{16*3}\] \[\Large 2i*\sqrt{6}+ (-1)*\sqrt{16}*\sqrt{3}\] \[\Large 2i*\sqrt{6}+ (-1)*4*\sqrt{3}\] \[\Large 2i*\sqrt{6}-4*\sqrt{3}\] \[\Large -4*\sqrt{3}+2i*\sqrt{6}\]

OpenStudy (anonymous):

you can also proceed as follows \[\sqrt{-6}=\sqrt{6}i\] and \[\sqrt{-8}=\sqrt{8}i\] maybe that would help then it is the distributive law recalling that when you see \(i^2\) you replace it by \(-1\)

OpenStudy (anonymous):

Nooow I got it, lol thank you Jim.

jimthompson5910 (jim_thompson5910):

You can optionally rearrange the terms to get \[\Large -4*\sqrt{3}+2\sqrt{6} * i\]